OpenStudy (austinl):

A uniformly charged conducting sphere of 0.93 m diameter has a surface charge density of 9.5 µC/m^2. (a) Find the net charge in coulombs on the sphere. (b) What is the total electric flux leaving the surface of the sphere?

OpenStudy (austinl):

For part (a) I got an answer of .000056 C... part (b) I am unsure.

OpenStudy (anonymous):

Again use Gauss's law

OpenStudy (anonymous):

"flux x area = q / epsilonsub0" @douglaswinslowcooper this is wrong :-/

OpenStudy (austinl):

For the first part, I did that. Except I forgot the squared. \(\Large{q=4\cdot\pi\cdot(.465)^2=2.71716}\) Then I just use, \(\Large{\Phi=\frac{q}{\epsilon_{0}}=\frac{2.71716}{8.85\times10^{-12}}=3.07024\times10^{11}}\) Of course, sig figs are off.

OpenStudy (anonymous):

They have asked total electric flux leaving the surface.. so u just some close gaussian surface enclosing the sphere and then gauss laws says the net flus through that surface would be q/eps so that must be your answer!

OpenStudy (anonymous):

yea sig figs are WAY off :D :D!!

OpenStudy (austinl):

So... did I actually manage to do it correctly?

OpenStudy (anonymous):

well.. you did :P

OpenStudy (austinl):

Awesome! :D Thanks guys!

OpenStudy (anonymous):

but what is important is whether you understood it or not!

OpenStudy (austinl):

Yeah, I am getting this stuff better and better :P

OpenStudy (anonymous):

kudos!!! :D

OpenStudy (austinl):

Wait.... I just plugged in these answers and it says they are wrong.

OpenStudy (austinl):

Here we are.

OpenStudy (austinl):

Wait, first one I needed to multiply by 9.5E-6

OpenStudy (austinl):

(a) .000026 (b) \(2.93785\times10^{6}\) Look better?

OpenStudy (anonymous):

q IS NOT EQUAL TO 4 pi R square... lolzz haha!! q = (density) times (4 pi R^2)

OpenStudy (austinl):

Both are correct now! :D

OpenStudy (austinl):

I forgot to multiply by density lmao.

OpenStudy (anonymous):

@Mashy Yes. I was wrong. Flux is integral of E field times area. Closed surface integral (Gauss) would = q/epsilonsub0. I will delete.

OpenStudy (anonymous):

@douglaswinslowcooper yup .. just a silly mistake.. but not as silly as austinL dude's mistake :D :D!!

OpenStudy (anonymous):

We are racing to the bottom!